# Multivariate de Bruijn Graphs: A Symbolic Graph Framework for Time Series Forecasting

### v1

@@ 1 Introduction
Foundation models have made great strides in NLP [15] and computer vision [1] by utilizing non-numeric, large-scale data; however, the same level of generalization has not emerged in the domain of time series where challenges exist e.g., temporal heterogeneity, multivariate dependencies, and distant relationships.
@@
Existing time series forecasting models rely heavily on the inputs being numeric, especially continuous [4] [10] while missing out on recurring motifs that are central to many real-world temporal processes. Moreover, the common practice of creating fixed-size sliding window fragments from long sequences to facilitate supervised training limits the model’s access to the global context outside the current fragment. To address these gaps, we introduce a new architecture that incorporates a symbolic graph-based structure into time series modeling through de Bruijn Graphs (dBG) [3]. dBGs are widely used in computational biology in applications such as genome assembly and k-mer–based (also known as k-tuple) sequence modeling to efficiently and compactly represent long sequences from overlapping fragments [2]. Modeling time series data as dBG allows transforming continuous temporal data into a compact symbolic representation that captures recurring motifs and long-range dependencies across sequence fragments. We introduce a novel method for encoding multivariate time series as Multivariate de Bruijn Graphs (MdBGs) to capture patterns both within and across dimensions, allowing models to access cross sequence global context while processing each sliding window locally.
Our contributions are threefold: (i) MdBG, a multivariate extension of de Bruijn graphs, captures intra-, inter-dimensional, and long-range dependencies; (ii) DRAGON, an encoder built on MdBGs, generates task-specific time series embeddings and integrates into any model; and (iii) DRAGON outperforms state-of-the-art methods on forecasting benchmarks, including TimesNet [13].
@@ 2.1.1 Data Processing
Let $\mathcal{D}^{\text{raw}}=\{\mathbf{X}_{1}^{\text{raw}},\mathbf{X}_{2}^{\text{raw}},\dots,\mathbf{X}_{D}^{\text{raw}}\}$ be a time series data with $D$ dimensions. Each raw sequence is defined as:
$\mathbf{X}_{i}^{\text{raw}}=[x_{1}^{i},x_{2}^{i},\dots,x_{S}^{i}],\quad$
where $x_{t}^{i}\in\mathbb{R}$ and $S$ is the sequence length. Since dBGs operate on categorical data, the raw input $\mathbf{X}_{i}^{\text{raw}}$ is first discretized using a discretization function $\textsc{DISC}_{i}(\mathbf{X}_{i}^{\text{raw}},\alpha_{i})\rightarrow\mathbf{X}_{i}^{\text{disc}}$. The set of bin sizes for all dimensions is denoted by $A=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$, where $\alpha_{i}$ represents the bin size for dimension $i$. Similarly, the set of discretization functions is denoted by $\mathbf{F}=\{\textsc{DISC}_{1},\textsc{DISC}_{2},\dots,\textsc{DISC}_{D}\}$. This formulation allows discretization to be customized independently for each dimension. The resulting discretized dataset, denoted as $\mathcal{D}$, consists of sequences defined as follows:
$\mathcal{D}^{\text{disc}}=\{\mathbf{X}_{1}^{\text{disc}},\mathbf{X}_{2}^{\text{disc}},\dots,\mathbf{X}_{D}^{\text{disc}}\},\quad\text{where}$
$\mathbf{X}_{i}^{\text{disc}}=[c_{1}^{i},c_{2}^{i},\dots,c_{S}^{i}],\quad c_{t}^{i}\in\{1,2,\dots,\alpha_{i}\}.$
Each discretized element $c_{t}^{i}$ represents a categorical value bounded by the discretization level $\alpha_{i}$.
@@ 2.1.2 Graph Construction
More formally, a dBG of order $k$ is a directed graph where each node represents a unique univariate $(k{-}1)$-tuple (a subsequence of length $k{-}1$), and each edge corresponds to a $k$-tuple (length-$k$ subsequence) observed in the input sequence. Specifically, the set of $k$-tuples.
$\mathbf{e}_{t}^{i}=\left(x_{t}^{i},x_{t+1}^{i},\dots,x_{t+k-1}^{i}\right)$
induces a directed edge from the prefix node to the suffix node,
$\mathbf{v}_{\text{prefix}}=\left(x_{t}^{i},\dots,x_{t+k-2}^{i}\right)\longrightarrow\mathbf{v}_{\text{suffix}}=\left(x_{t+1}^{i},\dots,x_{t+k-1}^{i}\right)$
for $t=1,2,\dots,\left(S-k+1\right)$. Each $\mathbf{e}_{t}^{i}$ defines a directed edge between two $(k{-}1)$-tuple nodes, capturing temporal transitions in the sequence. Repeated transitions increment the edge weights, encoding frequency. The construction of a dBG from univariate data is illustrated in Figure 1 (Left).
@@
Once MdBG is constructed from the training set, Graph Diffusion Convolution (GDC) is applied to further refine the structural properties of the MdBG and enhance information propagation across the constructed multivariate graph. GDC modifies the adjacency matrix by incorporating global diffusion patterns, effectively capturing long-range dependencies and smoothing over noisy or sparse connections. Specifically, we utilize Personalized PageRank (PPR)-based diffusion, which has demonstrated improved performance in node representation learning tasks by preserving both local and global graph structure [6]. This transformation improves the robustness and expressiveness of the learned node embeddings, which are later used in downstream tasks.
Table 1: Average MSE and MAE per model across four datasets (ETTh1, ETTh2, ETTm1, ETTm2). Lowest values in each column are highlighted in bold.
@@
Transformer [11] 0.909 0.733 2.54 1.219 0.889 0.672 1.404 0.844
TSMixer [8] 1.002 0.756 2.262 1.236 0.916 0.690 0.907 0.723
Crossformer [4] 0.691 0.609 1.982 1.060 0.718 0.588 1.005 0.675
FiLM [16] 0.712 0.550 0.450 0.433 1.031 0.626 0.350 0.375
Nonstationary [9] 0.742 0.586 0.576 0.506 0.764 0.557 0.418 0.399
PatchTST [10] 0.642 0.525 0.448 0.432 1.014 0.624 0.347 0.373
TimeMixer [12] 0.633 0.519 0.447 0.434 0.942 0.601 0.343 0.370
Autoformer [14] 0.656 0.556 0.458 0.450 0.854 0.603 0.335 0.376
TimeXer [8] 0.613 0.512 0.444 0.428 0.904 0.589 0.336 0.362
TimesNet [13] 0.639 0.525 0.457 0.438 0.850 0.578 0.332 0.358
@@ Appendix B MdBG Construction Algorithm
The MdBG construction algorithm is shown in Algorithm 1. The input consists of multivariate continous time series data $\mathcal{D}^{\text{raw}}=\{\textbf{X}_{1}^{\text{raw}},\dots,\textbf{X}_{m}^{\text{raw}}\}$, where each $\textbf{X}_{i}^{\text{raw}}$ denotes a univariate sequence. $k$ denotes the $k$-tuple length and the =$\mathcal{F}=\{\mathrm{DISC}_{1},\dots,\mathrm{DISC}_{m}\}$ represents the discretization functions for each dimension which can be shared or unique. First, an empty graph is initialized $G=(\mathcal{V},\mathcal{E})$ (line 2). For each input dimension $i$, the data in that dimension is discretized $\mathrm{DISC}_{i}$ (line 3-6). A sliding window of size $k$ is applied to the discretized data to generate $k$-tuples $R_{i}^{t}$ and $K_{i}^{t}$ (lines 9-10). Extracted $k$-tuple is used to generate a prefix $u_{i}$ and a suffix node $v_{i}$ and its feature set $F_{u_{i}}$ or $F_{v_{i}}$ is initialized (lines 11–20). These feature sets are always updated to track which raw sequences contributed to each node (lines 21–22). If the edge $(u_{i},v_{i})$ does not exist in $\mathcal{E}$, then it is initialized with the edge weight $w=1$, otherwise $w$ is incremented each time it is visited to reflect the frequency (lines 23–28). At the initial time step $t=0$, all prefix nodes $u_{i}$ across dimensions are connected using a bidirectional hyper-tuple edge to capture the inter-dimensional dependencies (lines 30–33). For all subsequent timesteps, hyper-tuple edges are formed between suffix nodes $v_{i}$ as well (lines 35–37). Finally, the constructed MdBG ($G$) is returned (line 39).
@@
1
Input :  $\mathcal{D}^{\text{raw}}=\{\textbf{X}_{1}^{\text{raw}},\textbf{X}_{2}^{\text{raw}},\dots,\textbf{X}_{m}^{\text{raw}}\}$: Aligned multivariate time series,
@@
2
Output :  $G=(\mathcal{V},\mathcal{E})$: Constructed Multivariate de Bruijn Graph
3
4
5Function MdBGConstruction($\mathcal{D}^{\text{raw}},k,\mathbf{F}$):
6       Initialize empty graph: $G\leftarrow(\mathcal{V}=\varnothing,\mathcal{E}=\varnothing)$
7
8      for $i\leftarrow 1$ in $m$ do
9             $\textsc{DISC}_{i}\leftarrow\mathbf{F}[i]$
10            $\textbf{X}_{i}^{\text{disc}}\leftarrow\textsc{DISC}_{i}(\textbf{X}^{\text{raw}}_{i})$
11
@@ 10            $\textbf{X}_{i}^{\text{disc}}\leftarrow\textsc{DISC}_{i}(\textbf{X}^{\text{raw}}_{i})$
16                   $K_{i}^{t}\leftarrow\textbf{X}_{i}^{\text{disc}}[t:t+k]$
17
18                  $R_{i}^{t}\leftarrow\textbf{X}_{i}^{\text{raw}}[t:t+k]$
19
20                  $u_{i}\leftarrow(i,K_{i}^{t}[0:k{-}1])$
21
22                  $v_{i}\leftarrow(i,K_{i}^{t}[1:k])$
23
24                  if $u_{i}\notin\mathcal{V}$ then
25                         Add $u_{i}$ to $\mathcal{V}$
26
27                        $\mathcal{F}_{u_{i}}=\varnothing$
@@
30                         Add $v_{i}$ to $\mathcal{V}$
31
32                        $\mathcal{F}_{v_{i}}=\varnothing$
@@
34
35                  $\mathcal{F}_{u_{i}}=\mathcal{F}_{u_{i}}\cup\{R_{i}^{t}[0:k{-}1]\}$
36                  $\mathcal{F}_{v_{i}}=\mathcal{F}_{v_{i}}\cup\{R_{i}^{t}[1:k]\}$
37                  if $(u_{i},v_{i})\notin\mathcal{E}$ then
@@ C.2 Extended Experimental Results
All experiments use an input sequence length of 12. We employ three DRAGON modules with order $k=4$ and discretization alphabet sizes $\alpha=20$, $25$, and $30$, respectively for all layers/dimensions. Similarly, a uniform discretization function is used, which assigns equal-width bins across all features and dimensions for every layer. Each DRAGON module incorporates two layers of graph attention, and Graph Diffusion Convolution (GDC) is applied with a top-$k$ value of 32. For the downstream TimesNet model, the “optimal” hyperparameters recommended for each benchmark dataset and forecasting horizon is adopted. The graph embedding dimension $G$ is consistently set equal to the model channel size $C$. Our experimental pipeline, along with the hyperparameter configurations for all models, follows the TSLib framework [13], available at: https://github.com/thuml/Time-Series-Library.
@@
The graph construction algorithm is implemented using the NetworkX library [7], and the resulting graphs are subsequently converted into Data objects compatible with PyTorch Geometric [5]. All experiments are conducted on a single NVIDIA A100 GPU with 42 GB of memory.

### v3

@@ 1 Introduction
Foundation models have made great strides in NLP (Zhao et al., 2023) and computer vision (Awais et al., 2023) by utilizing non-numeric, large-scale data; however, the same level of generalization has not emerged in the domain of time series where challenges exist e.g., temporal heterogeneity, multivariate dependencies, and distant relationships.
@@
Existing time series forecasting models rely heavily on the inputs being numeric, especially continuous (Ekambaram et al., 2024) (Nie et al., 2023) while missing out on recurring motifs that are central to many real-world temporal processes. Moreover, the common practice of creating fixed-size sliding window fragments from long sequences to facilitate supervised training limits the model’s access to the global context outside the current fragment. To address these gaps, we introduce a new architecture that incorporates a symbolic graph-based structure into time series modeling through de Bruijn Graphs (dBG) (De Bruijn, 1946). dBGs are widely used in computational biology in applications such as genome assembly and k-mer–based (also known as k-tuple) sequence modeling to efficiently and compactly represent long sequences from overlapping fragments (Compeau et al., 2011). Recently, dBGs have also been explored in the context of time series modeling for capturing symbolic temporal patterns and enhancing forecasting performance (Cakiroglu et al., 2024b, a). Modeling time series data as dBG allows transforming continuous temporal data into a compact symbolic representation that captures recurring motifs and long-range dependencies across sequence fragments. We introduce a novel method for encoding multivariate time series as Multivariate de Bruijn Graphs (MdBGs) to capture patterns both within and across dimensions, allowing models to access cross sequence global context while processing each sliding window locally.
Our contributions are threefold: (i) MdBG, a multivariate extension of de Bruijn graphs, captures intra-, inter-dimensional, and long-range dependencies; (ii) DRAGON, an encoder built on MdBGs, generates task-specific time series embeddings and integrates into any model; and (iii) DRAGON outperforms state-of-the-art methods on forecasting benchmarks, including TimesNet (Wu et al., 2022).
@@ 2.1.1 Data Processing
Let $\mathcal{D}^{\text{raw}}=\{\mathbf{X}_{1}^{\text{raw}},\mathbf{X}_{2}^{\text{raw}},\dots,\mathbf{X}_{D}^{\text{raw}}\}$ be a time series data with $D$ dimensions. Each raw sequence is defined as: $\mathbf{X}_{i}^{\text{raw}}=[x_{1}^{i},x_{2}^{i},\dots,x_{S}^{i}],$ where $x_{t}^{i}\in\mathbb{R}$ and $S$ is the sequence length. Since dBGs operate on categorical data, the raw input $\mathbf{X}_{i}^{\text{raw}}$ is first discretized using a discretization function $\textsc{DISC}_{i}(\mathbf{X}_{i}^{\text{raw}},\alpha_{i})\rightarrow\mathbf{X}_{i}^{\text{disc}}$. The set of bin sizes for all dimensions is denoted by $A=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$, where $\alpha_{i}$ represents the bin size for dimension $i$. Similarly, the set of discretization functions is denoted by $\mathbf{F}=\{\textsc{DISC}_{1},\textsc{DISC}_{2},\dots,\textsc{DISC}_{D}\}$. This formulation allows discretization to be customized independently for each dimension. The resulting discretized dataset, denoted as $\mathcal{D}$, consists of sequences defined as follows: $\mathcal{D}^{\text{disc}}=\{\mathbf{X}_{1}^{\text{disc}},\mathbf{X}_{2}^{\text{disc}},\dots,\mathbf{X}_{D}^{\text{disc}}\},$ where $\mathbf{X}_{i}^{\text{disc}}=[c_{1}^{i},c_{2}^{i},\dots,c_{S}^{i}],\quad c_{t}^{i}\in\{1,2,\dots,\alpha_{i}\}.$ Each discretized element $c_{t}^{i}$ represents a categorical value bounded by the discretization level $\alpha_{i}$.
@@ 2.1.2 Graph Construction
More formally, a dBG of order $k$ is a directed graph where each node represents a unique univariate $(k{-}1)$-tuple (a subsequence of length $k{-}1$), and each edge corresponds to a $k$-tuple (length-$k$ subsequence) observed in the input sequence. Specifically, the set of $k$-tuples $\mathbf{e}_{t}^{i}=\left(x_{t}^{i},x_{t+1}^{i},\dots,x_{t+k-1}^{i}\right)$ induces a directed edge from the prefix node to the suffix node, $\mathbf{v}_{\text{prefix}}=(x_{t}^{i},\dots,x_{t+k-2}^{i})\rightarrow\mathbf{v}_{\text{suffix}}=(x_{t+1}^{i},\dots,x_{t+k-1}^{i})$ for $t=1,2,\dots,\left(S-k+1\right)$. Each $\mathbf{e}_{t}^{i}$ defines a directed edge between two $(k{-}1)$-tuple nodes, capturing temporal transitions in the sequence. Repeated transitions increment the edge weights, encoding frequency. The construction of a dBG from univariate data is illustrated in Figure 1 (Left).
@@
Once MdBG is constructed from the training set, Graph Diffusion Convolution (GDC) is applied to further refine the structural properties of the MdBG and enhance information propagation across the constructed multivariate graph. GDC modifies the adjacency matrix by incorporating global diffusion patterns, effectively capturing long-range dependencies and smoothing over noisy or sparse connections. Specifically, we utilize Personalized PageRank (PPR)-based diffusion, which has demonstrated improved performance in node representation learning tasks by preserving both local and global graph structure (Gasteiger et al., 2022). This transformation improves the robustness and expressiveness of the learned node embeddings, which are later used in downstream tasks.
Table 1: Average MSE and MAE per model across four datasets (ETTh1, ETTh2, ETTm1, ETTm2). Lowest values in each column are highlighted in bold. The DRAGON module achieves the lowest error rates in the majority of settings, demonstrating its effectiveness in recovering long-range dependencies from severely limited input contexts. These results are obtained using a fixed input length of 12 and illustrate the generalization capabilities of DRAGON in multivariate time series benchmarks.
@@
Transformer (Vaswani et al., 2017) 0.909 0.733 2.54 1.219 0.889 0.672 1.404 0.844
TSMixer (Lin et al., 2024) 1.002 0.756 2.262 1.236 0.916 0.690 0.907 0.723
Crossformer (Ekambaram et al., 2024) 0.691 0.609 1.982 1.060 0.718 0.588 1.005 0.675
FiLM (Zhou et al., 2022) 0.712 0.550 0.450 0.433 1.031 0.626 0.350 0.375
Nonstationary (Liu et al., 2022) 0.742 0.586 0.576 0.506 0.764 0.557 0.418 0.399
PatchTST (Nie et al., 2023) 0.642 0.525 0.448 0.432 1.014 0.624 0.347 0.373
TimeMixer (Wang et al., 2024) 0.633 0.519 0.447 0.434 0.942 0.601 0.343 0.370
Autoformer (Wu et al., 2021) 0.656 0.556 0.458 0.450 0.854 0.603 0.335 0.376
TimeXer (Lin et al., 2024) 0.613 0.512 0.444 0.428 0.904 0.589 0.336 0.362
TimesNet (Wu et al., 2022) 0.639 0.525 0.457 0.438 0.850 0.578 0.332 0.358
@@ Appendix B MdBG Construction Algorithm
The MdBG construction algorithm with the time complexity of $\mathcal{O}(m^{2}\cdot T)$ is shown in Algorithm 1. The input consists of multivariate continous time series data $\mathcal{D}^{\text{raw}}=\{\textbf{X}_{1}^{\text{raw}},\dots,\textbf{X}_{m}^{\text{raw}}\}$, where each $\textbf{X}_{i}^{\text{raw}}$ denotes a univariate sequence. $k$ denotes the $k$-tuple length and the =$\mathcal{F}=\{\mathrm{DISC}_{1},\dots,\mathrm{DISC}_{m}\}$ represents the discretization functions for each dimension which can be shared or unique. First, an empty graph is initialized $G=(\mathcal{V},\mathcal{E})$ (line 2). For each input dimension $i$, the data in that dimension is discretized $\mathrm{DISC}_{i}$ (line 3-6). A sliding window of size $k$ is applied to the discretized data to generate $k$-tuples $R_{i}^{t}$ and $K_{i}^{t}$ (lines 9-10). Extracted $k$-tuple is used to generate a prefix $u_{i}$ and a suffix node $v_{i}$ and its feature set $F_{u_{i}}$ or $F_{v_{i}}$ is initialized (lines 11–20). These feature sets are always updated to track which raw sequences contributed to each node (lines 21–22). If the edge $(u_{i},v_{i})$ does not exist in $\mathcal{E}$, then it is initialized with the edge weight $w=1$, otherwise $w$ is incremented each time it is visited to reflect the frequency (lines 23–28). At the initial time step $t=0$, all prefix nodes $u_{i}$ across dimensions are connected using a bidirectional hyper-tuple edge to capture the inter-dimensional dependencies (lines 30–33). For all subsequent timesteps, hyper-tuple edges are formed between suffix nodes $v_{i}$ as well (lines 35–37). Finally, the constructed MdBG ($G$) is returned (line 39).
@@
Algorithm 1 MdBG Construction Algorithm
Input :
$\mathcal{D}^{\text{raw}}=\{\textbf{X}_{1}^{\text{raw}},\textbf{X}_{2}^{\text{raw}},\dots,\textbf{X}_{m}^{\text{raw}}\}$: Aligned multivariate time series,
@@
Output :
$G=(\mathcal{V},\mathcal{E})$: Constructed Multivariate de Bruijn Graph
Function MdBGConstruction($\mathcal{D}^{\text{raw}},k,\mathbf{F}$):
2       Initialize empty graph: $G\leftarrow(\mathcal{V}=\varnothing,\mathcal{E}=\varnothing)$for $i\leftarrow 1$ in $m$ do
3             $\textsc{DISC}_{i}\leftarrow\mathbf{F}[i]$ $\textbf{X}_{i}^{\text{disc}}\leftarrow\textsc{DISC}_{i}(\textbf{X}^{\text{raw}}_{i})$
@@
7                   $K_{i}^{t}\leftarrow\textbf{X}_{i}^{\text{disc}}[t:t+k]$$R_{i}^{t}\leftarrow\textbf{X}_{i}^{\text{raw}}[t:t+k]$$u_{i}\leftarrow(i,K_{i}^{t}[0:k{-}1])$$v_{i}\leftarrow(i,K_{i}^{t}[1:k])$if $u_{i}\notin\mathcal{V}$ then
8                         Add $u_{i}$ to $\mathcal{V}$$\mathcal{F}_{u_{i}}=\varnothing$
@@
11                         Add $v_{i}$ to $\mathcal{V}$$\mathcal{F}_{v_{i}}=\varnothing$
@@
13                  $\mathcal{F}_{u_{i}}=\mathcal{F}_{u_{i}}\cup\{R_{i}^{t}[0:k{-}1]\}$ $\mathcal{F}_{v_{i}}=\mathcal{F}_{v_{i}}\cup\{R_{i}^{t}[1:k]\}$ if $(u_{i},v_{i})\notin\mathcal{E}$ then
@@ C.2 Extended Experimental Results
All experiments use an input sequence length of 12. We employ three DRAGON modules with order $k=4$ and discretization alphabet sizes $\alpha=20$, $25$, and $30$, respectively for all layers/dimensions. Similarly, a uniform discretization function is used, which assigns equal-width bins across all features and dimensions for every layer. Each DRAGON module incorporates two layers of graph attention, and Graph Diffusion Convolution (GDC) is applied with a top-$k$ value of 32. For the downstream TimesNet model, the “optimal” hyperparameters recommended for each benchmark dataset and forecasting horizon is adopted. The graph embedding dimension $G$ is consistently set equal to the model channel size $C$. Our experimental pipeline, along with the hyperparameter configurations for all models, follows the TSLib framework (Wu et al., 2022), available at: https://github.com/thuml/Time-Series-Library.
@@
The graph construction algorithm is implemented using the NetworkX library (Hagberg et al., 2008), and the resulting graphs are subsequently converted to data objects compatible with PyTorch Geometric (Fey & Lenssen, 2019). All experiments are conducted on a single NVIDIA A100 GPU with 42 GB of memory.
@@
C.3 Future Work
We plan to extend the DRAGON framework in several promising directions. The first step is to improve its scalability in terms of graph construction and masking strategies to better accommodate long-term forecasting and high dimensional time series data. Subsequently, we aim to conduct more extensive experiments across diverse benchmark datasets from various domains and multiple downstream models to assess generalization. Beyond forecasting, we intend to transform DRAGON into a multimodal architecture by incorporating support for additional downstream tasks, including anomaly detection, classification, and imputation. Ultimately, our long-term goal is to evolve DRAGON into a foundation model by training it on large-scale, heterogeneous time series data in a self-supervised manner.